# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### v2

@@ 4 MrDeepFakes Tube Site
Founded in 2018, MrDeepFakes experienced relatively little engagement during its first few years online [timmerman2023studying]. However, following a broader spike in sexual deepfake interest starting in 2021 (Figure 2(a)), MrDeepFakes’ popularity has taken off. As The site adopted an interface that emphasized watching uploaded sexual deepfake videos, and as of July 2024, MrDeepFakes is one of the 10K most popular websites globally according to Google CrUX [ruth2022toppling][durumeric2023cached, ruth2022toppling]. At the time of our data collection in November 2023, the site hosts hosted $42986$ sexual deepfake videos posted by 1,880 users, which have garnered over 1.5B views and a median 14.3K views per video. In this section, we analyze the individuals targeted by deepfake videos uploaded to MrDeepFakes .
@@ 4.1 Methodology
To understand the individuals targeted by deepfakes, we analyzed uploaded videos on the MrDeepFakes tube site. Posted deepfake videos have attached metadata that enables discoverability on the site, including the name of the depicted (i.e., deepfaked) target, descriptive categories (Table 3), and free-form tags. Over 90% of videos specified the depicted celebrities, which enabled us to identify and analyze the demographics of 3,803 unique individuals depicted in deepfake videos.
@@ 4.2 Targets of Deepfake Videos
Posted deepfake videos have attached metadata that enables discoverability on the site, including the name of the depicted (i.e., deepfaked) target, descriptive categories (Table 3), and free-form tags. Over 90% of videos indicated the depicted celebrities, which enabled us to identify and analyze the demographics of 3,803 unique individuals depicted in deepfake videos.
Targeted individuals follow a long-tail distribution: the ten most targeted celebrities are depicted in 400–900 videos (1–2% of all posted videos) each, but over 40% of targets are depicted in only a single one video. We manually investigated the $1942$ celebrities most commonly depicted, who account for over 95% of videos, on the basis of gender/sexual identity, occupation, nationality, and social media following using Google search, news stories, celebrity websites, and public social media profiles.
The 20 most targeted individuals are all well-known women actresses, or popular American or Korean musicians. Just over one third of targeted individuals are American, followed by Korean individuals Koreans (11%), many of whom are Korean pop (“KpopK-pop”) celebrities. MrDeepFakes’ focus on sexual deepfakes rather than political disinformation appears to minimize the number of depicted geopolitical figures (43 individuals, 2.21%).
Consistent with industry reports [homesecurityheroes2023state, deeptrace2019state], we find that 91% of videos target gender-marginalized individuals (i.e., anyone whose labels do not consist solely of “man” or “unknown,” including women and LGBTQIA+ people). Of the $1942$ targeted individuals, 95.3% are women, and 2.1% identify as LGBTQIA+. Most targeted individuals are actors, musicians, television personalities, or social media influencers (Table 2). Actors comprise more than half of the targets (58.19%), possibly in part because there is more high quality more high-quality and readily available facial data exists in the form of public images and videos.
Top Video Categories—The most popular deepfake video categories. Categories are not mutually exclusive.
@@ 4.3 Non-Celebrity Targets
MrDeepFakes advertises itself as a “celebrity deepfake porn tube site,” and its community codifies accepting only sexual deepfakes that depict celebrities. Website rules state, among other requirements, that a celebrity must have “significant influence” on social media with over 200K followers on TikTok or 100K followers on Instagram, YouTube, Twitter, or Twitch, or 200K followers on TikTok. Despite these rules, 271 of the 1,942 individuals we investigated (14%) fall far below these thresholds. For 29 targets, we found no online presence, and for another 242 targets, we could not find profiles on any of the listed social media platforms that satisfied the minimum following criteria.
In one exampleof non-celebrity video uploads, we find that 38 Guatemalan newscasters with little to no social media following appear in over 300 deepfake videos. All of these videos were posted by two users, who both describe their focus on Guatemalan and Latin American individuals in their profiles.
These statistics likely significantly underestimate how often non-celebrities are targeted in sexual deepfake videos because our investigation is limited to the top 50% of the most commonly depicted individuals in public videos. At the same time, we also acknowledge the possibility of targets deleting social media accounts after, or perhaps due to, being depicted in deepfake content. Therefore, our measurements may also potentially overestimate non-celebrity targets.
@@ 4.4 Assault- and Abuse-Based Content
MrDeepFakes rules prohibit posting of content that “depicts celebrities in a negative manner.” While the rules do not define “negative manner,” is not defined, they list “rape scenes, abuse, [and] humiliation” as prohibited examples, noting that users who post such content will receive a warning if the source video is “professional porn” and will otherwise be banned. Despite thisprohibition, searches for terms related to violence, assault, rape, abuse, humiliation, and “gang bang” yield 1,017 videos (2.37%). Some depict the targeted individual as the perpetrator, rather than the victim, of such abuse, going beyond nonconsensually sexualizing targets to creating slanderous and criminal imagery.
According to the site footer, MrDeepFakes ’ stated purpose is to “provide a safe-haven without censorship” for people to experiment with deepfake technology, within the rules of the community. While this “safe-haven” has welcomed a substantial influx of new members, our findings indicate that deepfake creators are undeterred from generating harmful and inherently nonconsensual content and publicly distributing it on MrDeepFakes , violating even the code of ethics expressly defined by their own community.
@@ 5 MrDeepFakes Marketplace
In addition to its tube site, MrDeepFakes hosts a forum-based marketplace for members to request free or paid sexual deepfake videos. Aligned with the site’s overall growth, public requests for deepfake videos more than tripled since 2021 (Figure 2(b)). While there are also forum sections dedicated to the paid requests of celebrity and sex worker image datasets and trained models, they appear not to be used, with only three threads requesting celebrity datasets and none in the other two. In this section, we analyze requests for deepfake videos and their responses to understand the content that buyers are looking for sought by buyers and the costs associated with purchasing custom sexual deepfake content.
@@ 5.1 Methodology
We perform thematic analysis [braun2012thematic, kiger2020thematic] for a random subset of 263 (of 830) paid requests for deepfake videos. Our initial codes were guided by primary topics of interest from manual exploration of the marketplace and related work (e.g., buyer motivations grounded in those codified by Okolie et al. [okolie2023ibsa] for IBSA perpetrators, buyers’ methods for vetting sellers, desirable attributes of deepfake videos). Codes were iteratively added to the codebook (Appendix E) as we processed our dataset and adjusted recursively as our process surfaced themes (e.g., the degree to which buyers and sellers positioned themselves as artistic contributors). Two researchers independently assigned codes to the posts, and codes were not mutually exclusive. Inter-rater reliability (IRR) was reasonably high with a Kupper-Hafner metric [kupperhafner1989irr] of 0.72. For posts where the set of assigned codes were not identical, the researchers met to discuss the conflicts and reach agreement to obtain a final coded dataset. Presented quotes were jointly paraphrased by two authors.
@@ 5.2 Requests for Deepfakes
The MrDeepFakes marketplace is request-based: buyers request free or commission paid deepfake videos in the Forum Section: Adult DeepFake Videos (NSFW/18+) / Video Requests. As of November 2023, there were 830 (66.9%) paid requests and 411 (33.1%) unpaid requests. 94.7% of requests were responded to: paid requests received an average 2.6 replies, while unpaid requests received an average of 4.6 replies. Over 90% of the 707 unique buyers made exactly one public request; this may be due to later transactions occurring privately once a buyer has identified a satisfactory seller.
We thematically analyzed a random subset of 263 requests out of the full set of 830 paid requests for deepfake videos. Our initial labels were guided by the primary topics of interest (e.g., buyer motivations grounded in those codified by Okolie et al. [okolie2023ibsa] for IBSA perpetrators, vetting sellers, desirable attributes of deepfake videos), and labels were iteratively added to the codebook (Appendix E) when new themes emerged. Two researchers independently assigned labels to the posts, and inter-rater reliability (IRR) was reasonably high with a Kupper-Hafner metric [kupperhafner1989irr] of 0.72. Both researchers met to resolve conflicts and obtain a final coded dataset.
The vast majority (96.2%) of paid requests on MrDeepFakes are for “adult deepfake” videos, as designated by its parent section (Appendix A). However, we also find requests for non-sexual or “SFW” deepfakes (4), machine learning models for creating deepfakes (3), and data to be used in deepfake creation (2). Seven buyers specified that making a deepfake request is what drove them to join MrDeepFakes or to post for the first time; 13 posts buyers requested multiple videos or promised longer-term patronage if satisfied with the finished product.
Requests provided a range of details for technical and target specification. 10.9% of requests are intentionally vague about the identity of their target. On the other handmany, m hand, many requests not only specify the their target, but they also discuss the availability of targets’ facial image data (e.g., linking to targets’ social media accounts, listing movies featuring the target). Further, 31 posts (11.7%) advertise that they can facilitate the creation process by providing creators access to private or personally-curated repositories of source or destination image data (e.g., high-resolution video files, personal videos taken of musician targets at shows).
@@ 5.2.1 Buyer Motivations
Prior work has predominantly described the motives for deepfake abuse as either seeking nonconsensual sexual gratification or producing disinformation [okolie2023ibsa, thomas2021hatesok, dhsincreasing]. While we do find sexual gratification to be a major motivator, we find others as well.
Nonconsensual Sexual Gratification.
Nonconsensual Sexual Gratification.Nearly a quarter (23.7%) of deepfake requests are explicitly connected to the buyer’s pleasure (e.g., sexual remarks about the target in the request post). Several posts specifically seek extreme (i.e., racialized or violent) or fetish content. [^3] We use the term “fetish” as defined in DSM-5, where someone’s sexual pleasure depends on nonliving objects (such as undergarments or high-heeled shoes) or a highly specific focus on a body part (most often nongenital) [kafka2010fetish]. The majority of the targets on MrDeepFakes are not personal acquaintances of the buyers or sellers of the content that depicts them. Therefore, the buyers’ motives are not captured by understandings of other IBSA perpetrators, such as producers of “revenge porn.” However, we do find instances of requests that demonstrate the parasocial nature of sexual deepfake contentdeepfake NCII. For example, in two requests for content depicting an actress and a social media influencer, buyers expressed perceived intimacy with their targets (e.g., referring to the target as “their” target, claiming that they not only “love” her, but that they love her “the way she is”).
Power and Masculinity.
@@
ID LOVE HER [T]O BE DONE FOR TWO REASONS, 1. she’s hot, 2. Shes a proud actor who has a huge ego, I think if this happened it might humble her and give her something to laugh about.
@@
[redacted], [stage name] now that shes doing porn but her looks have gone, she’d probably like having her old face reused for hard porn. IF you do I suggest maybe giving her a cut on mrdeepfakes, I think she’s strapped for cash.
@@
Likewise, one buyer clearly expresses violent attitudes toward women, stating that the creator would get “bonus points” if they can make it “rough” or “painful,” tying the buyer’s sexual gratification to violence against women. Another request infantilizes its target as a “daddy’s daughter,” using her depiction in sexual scenes to fantasize about physically punishing her. All of these buyers received replies from interested sellers, and none were reprimanded about the kind of content they requested. Psychology literature has investigated the norms of masculinity, finding that controlling and subordinating women are significant indicators of conformity to masculine norms [mahalik2003cmni]. The normalization of buyers’ hyper-masculine expressions and preferences for sexual content by other buyers and sellers demonstrates MrDeepFakes market participants’ conformity to prototypical masculinity.
Community Contribution.
Community Contribution.In 7% of requests, the buyer situates their request as benefiting the community, or even more broadly, consumers of pornography. The most common manifestation of this sentiment is the buyer expressing that they (1) failed to find content fitting their parameters on the site , and (2) are seeking to fill this “gap,” believing it would benefit the community, whose members share values and sexual preferences with the buyer. For example, one request appeals to potential creators by stating that their request would be a good fit for members of MrDeepFakes:
Let me know if anyone is interested, since I know A LOT of people are looking for something with her…Would be a good fit for Mr.DeepFakes and guarantee would be a very popular video since nothing else exists of her online outside Instagram/Twitch.
@@
I would be proud to commission the first digital masterpiece featuring Elizabeth Mitchell’s (not) breach into the hardcore and hope to find an artist willing to create such a work…I’m not in the position to pay top dollar so I hope to find a creator who shares my passion for this project.
@@
allow me to explain: it will seem strange to you that i request a deepfake between pornstars but for connoisseurs like me of the world of porn we know that [redacted] never did a BBC scene but her similar [redacted] did, so my request…
@@
The security community has taxonomized the harm of online abuse, characterizing perpetrators as driven by the desire to inflict physical, psychological, or sexual harm, silence, or coerce targets [thomas2021hatesok]. However, the perception of deepfakes as artand of their consumers as connoisseurs and of their consumers as connoisseurs introduces a new intent, which we discuss in Section 7.1. We study the deepfake creation process and how the MrDeepFakes community supports inexperienced creators in Section 6.
@@ 5.2.2 Requests for Non-Celebrity Deepfakes
Despite forum rules prohibiting requests for non-celebrity targets and requiring requests to include inclusion of target celebrities’ names (“[n]on-celebsare not allowed to be requested”” cannot be requested), 11% of requests did not identify their target. Furthermore, eight of these requests specified that the targets were not celebrities. Two in particular described targets from their the requesters’ personal lives: (1) a girl that the buyer “liked” and (2) the buyer’s partner.
These requests were not taken down or admonished by moderators, as we observed in another request thread where the buyer claimed to want a deepfake of themselves themselves and would “provide a (sic) consent for sure.” In this thread, a moderator, who conceded that MrDeepFakes did not have rules against “requesting deepfakes of yourself,” warned the buyer:
Keep in mind some users will want very solid evidence it’s actually you before they start working on it. Also judge whether you trust given creator, there is no guarantee [the video] won’t end up on the internet (not here, we don’t tolerate and don’t approve uploads of anything involving non-celebs, even if you upload it yourself, it’s celebs only here).
@@
[R]equest completed. Much appreciated. Top notch quality from a few photos and deepfake received in a week time. Would recommend him for anyone needing good fakes. Thanks for sending a sample before continuing with the final art allowing me to choose parts.
@@
The ability of buyers to find less scrupulous sellers parallels what prior investigations of underground markets have documented with alternative service providers who are tolerant of criminal or abusive behaviors [noroozian19platforms, kanich2008spamalytics, levchenko2011trajectories, motoyama11analysis, franklin2007inquiry, han2022infra]. The ambiguous and informal processes for verifying the alleged consent of private targets raise further concerns: buyers can abuse these gaps to thinly veil nonconsensual deepfake requests for any target as consensual and find willing deepfake creators. It is in these cases that we may anticipate the abuse of the paid requests section to create deepfakes of personal acquaintances or enemies of the buyer, which more closely resembles the motives underlying other forms of IBSA, such as sextortion, revenge, and reputational harm.

### v3

@@ 4 MrDeepFakes Tube Site
Founded in 2018, MrDeepFakes experienced relatively little engagement during its first few years online [62]. However, following a broader spike in sexual deepfake interest starting in 2021 (Figure 2(a)), MrDeepFakes’ popularity has grown rapidly. The site homepage’s interface emphasizes watching uploaded sexual deepfake videos, and as of July 2024, MrDeepFakes is one of the 10K most popular websites globally according to Google CrUX [12, 57]. At the time of our data collection in November 2023, the site hosted $42986$ sexual deepfake videos posted by 1,880 users, which garnered over 1.5B views and a median 14.3K views per video. In this section, we analyze the individuals targeted by deepfake videos uploaded to MrDeepFakes.
@@ 4.1 Methodology
To understand the individuals targeted by deepfakes, we analyzed videos uploaded to the MrDeepFakes tube site. Posted deepfake videos have attached metadata that enables discoverability on the site, including the name of the depicted (i.e., deepfaked) target, descriptive categories (Table 3), and free-form tags. Over 90% of videos specified the depicted celebrities, which enabled us to identify and analyze the demographics of 3,803 unique individuals depicted in deepfake videos.
@@ 4.2 Targets of Deepfake Videos
Targeted individuals follow a long-tailed distribution: the ten most targeted celebrities are depicted in 400–900 videos (1–2% of all videos) each, but over 40% of targets are depicted in only one video. We manually investigated the $1942$ celebrities most commonly depicted, who account for over 95% of videos, on the basis of gender/sexual identity, occupation, nationality, and social media following using Google search, news stories, celebrity websites, and social media profiles.
The 20 most targeted individuals are well-known women actresses, or popular American or Korean musicians. Just over one third of targeted individuals are American, followed by Koreans (11%), many of whom are Korean pop (“K-pop”) celebrities. MrDeepFakes’ focus on sexual rather than political disinformation deepfake content appears to minimize the number of depicted geopolitical figures (43 individuals, 2.21%).
Consistent with industry reports [20, 10], we find that 91% of videos target gender-marginalized individuals (i.e., anyone whose labels do not consist solely of “man” or “unknown,” including women and LGBTQIA+ people). Of the $1942$ targeted individuals, 95.3% are women, and 2.1% identify as LGBTQIA+. Most targeted individuals are actors, musicians, television personalities, or social media influencers (Table 2). Actors comprise more than half of the targets (58.19%), possibly in part because more high-quality and readily available facial data exists in the form of public images and videos.
@@ 4.3 Non-Celebrity Targets
MrDeepFakes advertises itself as a “celebrity deepfake porn tube site,” and its community codifies permitting only sexual deepfakes that depict celebrities. Website rules state, among other requirements, that a celebrity must have “significant influence” on social media with over 200K followers on TikTok or 100K followers on Instagram, YouTube, Twitter, or Twitch. Despite these rules, 271 of the 1,942 individuals we investigated (14%) fall far below these thresholds. For 29 targets, we found no online presence, and for another 242 targets, we could not find profiles on any of the listed social media platforms that satisfied the minimum following criteria. In one example, we find that 38 Guatemalan newscasters with little to no social media following appear in over 300 videos. All of these videos were posted by two users, who both describe specializing in Latin American individuals in their profiles.
These statistics likely significantly underestimate how often non-celebrities are targeted in sexual deepfake videos because our investigation is limited to the top 50% of the most commonly depicted individuals in public videos. At the same time, we acknowledge the possibility of targets deleting social media accounts after, or perhaps due to, being depicted in deepfake content. Therefore, our measurements may also potentially overestimate non-celebrity targets.
@@ 4.4 Assault- and Abuse-Based Content
MrDeepFakes rules prohibit posting of content that “depicts celebrities in a negative manner.” While “negative manner,” is not defined, they list “rape scenes, abuse, [and] humiliation” as prohibited examples, noting that users who post such content will receive a warning if the source video is “professional porn” and will otherwise be banned. Despite this, searches for terms related to violence, assault, rape, abuse, humiliation, and “gang bang” yield 1,017 videos (2.37%). Some depict the targeted individual as the perpetrator, rather than the victim, of such abuse, extending beyond nonconsensually sexualizing targets to creating slanderous and criminal imagery.
According to the site footer, MrDeepFakes’ stated purpose is to “provide a safe-haven without censorship” for people to experiment with deepfake technology, within the rules of the community. While this “safe-haven” has welcomed a substantial influx of new members, our findings indicate that deepfake creators are undeterred from generating harmful and inherently nonconsensual content and publicly distributing it on MrDeepFakes. Such behaviors violate even the code of ethics expressly defined by their own community.
@@ 5 MrDeepFakes Marketplace
In addition to its tube site, MrDeepFakes hosts a forum-based marketplace for members to request free or paid sexual deepfake videos. Aligned with the site’s overall growth, public requests for deepfake videos more than tripled since 2021 (Figure 2(b)). While there are also forum sections dedicated to the paid requests of celebrity and sex worker image datasets and trained models, they appear unused, with only three threads requesting celebrity datasets and none in the other two. In this section, we analyze requests for deepfake videos and their responses to understand the content sought by buyers and the costs associated with purchasing custom sexual deepfake content.
@@ 5.1 Methodology
We perform thematic analysis [5, 24] for a random subset of 263 (of 830) paid requests for deepfake videos. Our initial codes were guided by primary topics of interest from manual exploration of the marketplace and related work (e.g., buyer motivations grounded in those codified by Okolie et al. [45] for IBSA perpetrators, buyers’ methods for vetting sellers, desirable attributes of deepfake videos). Codes were iteratively added to the codebook (Appendix E) as we processed our dataset and adjusted recursively as our process surfaced themes (e.g., the degree to which buyers and sellers positioned themselves as artistic contributors). Two researchers independently assigned codes to the posts, and codes were not mutually exclusive. Inter-rater reliability (IRR) was reasonably high with a Kupper-Hafner metric [29] of 0.72. For posts where the set of assigned codes were not identical, the researchers met to discuss the conflicts and reach agreement to obtain a final coded dataset. Presented quotes were jointly paraphrased by two authors.
@@ 5.2 Requests for Deepfakes
The MrDeepFakes marketplace is request-based: buyers request free or commission paid deepfake videos in the forum section Adult DeepFake Videos (NSFW/18+) / Video Requests. As of November 2023, there were 830 (66.9%) paid and 411 (33.1%) unpaid requests. 94.7% of requests were responded to: paid requests received an average 2.6 replies, while unpaid requests received an average 4.6 replies. Over 90% of the 707 unique buyers made exactly one public request; this may be due to later transactions occurring privately once a buyer has identified a satisfactory seller.
The vast majority (96.2%) of paid requests on MrDeepFakes are for “adult deepfake” videos, as designated by its parent section (Appendix A). However, we also find requests for non-sexual or “SFW” deepfakes (4), machine learning models for creating deepfakes (3), and data used in deepfake creation (2). Seven buyers specified that making a deepfake request is what drove them to join MrDeepFakes or to post for the first time; 13 buyers requested multiple videos or promised longer-term patronage if satisfied with the finished product.
Requests provided a range of details for technical and target specification. 10.9% of requests are intentionally vague about the identity of their target. On the other handmany, m requests not only specify their target, but they also discuss the availability of targets’ facial image data (e.g., linking to targets’ social media accounts, listing movies featuring the target). Further, 31 posts (11.7%) advertise that they can facilitate the creation process by providing creators access to private or personally-curated repositories of source or destination image data (e.g., high-resolution video files, personal videos taken of musician targets at shows).
@@ 5.2.1 Buyer Motivations
Prior work has predominantly described the motives for deepfake abuse as either seeking nonconsensual sexual gratification or producing political disinformation [45, 61, 44]. While we do find sexual gratification to be a major motivator, we find others as well.
Nonconsensual Sexual Gratification. Nearly a quarter (23.7%) of requests are explicitly connected to the buyer’s pleasure (e.g., sexual remarks about the target in the request post). Several posts specifically seek extreme (i.e., racialized or violent) or fetish content. [^3] We use the term “fetish” as defined in DSM-5, where someone’s sexual pleasure depends on nonliving objects (such as undergarments or high-heeled shoes) or a highly specific focus on a body part (most often nongenital) [22]. The majority of the targets on MrDeepFakes are not personal acquaintances of the buyers or sellers of the content that depicts them. Therefore, the buyers’ motives are not captured by understandings of other IBSA perpetrators, such as producers of “revenge porn.” However, we do find instances of requests that demonstrate the parasocial nature of deepfake NCII. For example, in two requests for content depicting an actress and a social media influencer, buyers expressed perceived intimacy with their targets (e.g., referring to the target as “their” target, claiming that they not only “love” her, but that they love her “the way she is”).
@@
Likewise, one buyer clearly expresses violent attitudes toward women, stating that the creator would get “bonus points” if they can make it “rough” or “painful,” tying the buyer’s sexual gratification to violence against women. Another request infantilizes its target as a “daddy’s daughter,” using her depiction in sexual scenes to fantasize about physically punishing her. All of these buyers received replies from interested sellers, and none were reprimanded about the kind of content they requested. Psychology literature has investigated the norms of masculinity, finding that controlling and subordinating women are significant indicators of conformity to masculine norms [35]. The normalization of buyers’ hyper-masculine expressions and preferences for sexual content by other buyers and sellers demonstrates MrDeepFakes market participants’ conformity to prototypical masculinity.
Community Contribution. In 7% of requests, the buyer situates their request as benefiting the community, or even more broadly, consumers of pornography. The most common manifestation of this sentiment is the buyer expressing that they (1) failed to find content fitting their parameters on the site and (2) are seeking to fill this “gap,” believing it would benefit the community, whose members share values and sexual preferences with the buyer. For example, one request appeals to potential creators by stating that their request would be a good fit for members of MrDeepFakes:
@@
The security community has taxonomized the harms of online abuse, characterizing perpetrators as driven by the desire to inflict physical, psychological, or sexual harm to, silence, or coerce targets [61]. However, the perception of deepfakes as art and of their consumers as connoisseurs introduces a new intent, which we discuss in Section 7.1. We study the deepfake creation process and how the MrDeepFakes community supports inexperienced creators in Section 6.
@@ 5.2.2 Requests for Non-Celebrity Deepfakes
Despite forum rules prohibiting requests for non-celebrity targets and requiring inclusion of target celebrities’ names (“[n]on-celebs” cannot be requested), 11% of requests did not identify their target. Furthermore, eight of these requests specified that the targets were not celebrities. Two in particular described targets from the requesters’ personal lives: (1) a girl that the buyer “liked” and (2) the buyer’s partner.
These requests were not taken down or admonished by moderators, as we observed in another request thread where the buyer claimed to want a deepfake of themselves and would “provide a (sic) consent for sure.” In this thread, a moderator, who conceded that MrDeepFakes did not have rules against “requesting deepfakes of yourself,” warned the buyer:
@@
The ability of buyers to find less scrupulous sellers parallels what prior investigations of underground markets have documented with alternative service providers who are tolerant of criminal or abusive behaviors [43, 23, 32, 41, 14, 17]. The ambiguous and informal processes for verifying the alleged consent of private targets raise further concerns: buyers can abuse these gaps to thinly veil nonconsensual deepfake requests for any target as consensual and find willing deepfake creators. It is in these cases that we may anticipate the abuse of the paid requests marketplace to create deepfakes of personal acquaintances or enemies of the buyer, which more closely resembles the motives underlying other forms of IBSA, such as sextortion, revenge, and reputational harm.